bdrammeh:

a light bulb uses 40.0 Wof power when connected to a 110 V (DC) outlet. if the same bulb is connected to a european 220 V outlet, how much power will the bulb use? (hint: what is the resistance of the bulb?)(unit=W)

TheSmartOne:

Hi there. Do you have any idea which formula we should use? Hint: we want something that can help us calculate the power and we have the voltage and resistance (amps)

bdrammeh:

idk the formula

TheSmartOne:

Have you seen this formula before? Power = voltage * current P = IV Where power is in watts Current is A And voltage is V

bdrammeh:

P= IV=I^2R=v^2/R

bdrammeh:

AND YES IVE SEEN THAT FORMULA

TheSmartOne:

Yes, but we only need to select one of them :)

TheSmartOne:

PERFECTO Let's solve this It's ez pz

TheSmartOne:

So we know a light bulb uses 40.0 W of power and is connected to 110 V outlet Can you plug that into the formula and calculate the resistance of the bulb?

bdrammeh:

just do it for me please

bdrammeh:

so wat is the answer

TheSmartOne:

Unfortunately I can't do it for you, that would defeat the whole purpose of you being able to learn It'll take you less than two minutes to solve it if you work with me

bdrammeh:

smt

TheSmartOne:

P = I * V They gave us P = 40.0 W And V = 110 V Can you solve for I?

bdrammeh:

=302.5

TheSmartOne:

How'd you get that mhmm

bdrammeh:

its wrong

bdrammeh:

when a certain amount of current flows through a resistor, it uses 3.00 w of power. if the current doubles, how much power will the resistor use? (Hint: R stays the same. how does P depend on I and R?)(unit=W)

bdrammeh:

wat abou this one

TheSmartOne:

Did you find the correct answer then? 40 W = I * 110V Solve for I The units will be A Once you find I, you just multiply by 220 V to get the power

bdrammeh:

=66550

bdrammeh:

?

TheSmartOne:

uhhh wait Shouldn't we be using the P= V^2/R formula I accidentally confused current with resistance

TheSmartOne:

40 = (110)^2 / R Solve for R R = (110)^2 / 40 R = ?

bdrammeh:

302.5

bdrammeh:

the answer is 0.852

TheSmartOne:

How did you get 0.852? 302.5 ohms is the resistance Now for what we're trying to find P = V^2/R We know that V = 220 And R = 302.5 So what is P=?

bdrammeh:

=160

TheSmartOne:

And that's your answer 160 W